OpenStudy (anonymous):

ten inches of heavy, wet snow are equivalent to three inches of rain. Estimate the water content in 17 inches of heavy, wet snow.

OpenStudy (texaschic101):

Estimate ... 10 in snow to 3 in rain...20 in snow to 6 in....so my estimate would be a little less then 6 inches. actual answer ... 10/3 = 17/x (10 in snow to 3 in rain = 17 in snow to x in rain)cross multiply 10x = (17)(3) 10x = 51 10x/10 = 51/10 x = 5.1 in of rain. check... 10/3 = 17/5.1 (cross multiply) (10)(5.1) = (3)(17) 51 = 51 (correct)